2. Inputs and outputs Raphaels Performance Pizza is a small restaurant in San Diego that sells gluten-free pizzas. RaphaelsProduction Function QUANTITY OF OUTPUT (Pizzas) AHHO LABOR (Number of workers) Suppose that labor is Raphaels only variableTotal Cost TOTAL COST (Dollars) HAHAHA 0 + + + 0 20 180 200 40 60 80 100 120 140 160 QUANTITY OF OUTPUT (Pizzas) The law of d
Answer #1

Labor (L) Output MPL FC VC=30* L TC=FC+VC 50 60 90 120 120 140 150 150 120 150 Quantity of Output (Pizzas) 10 30 180 5 6 2 3

Law of diminishing marginal product is demonstrated by the fact that the total output increases at decreasing rate as more and more unit of labors are employed.